# AutoManual: Generating Instruction Manuals by LLM Agents via Interactive Environmental Learning

### v1

@@ Abstract
Large Language Models (LLM) based agents have shown promise in autonomously completing tasks across various domains, e.g., robotics, games, and web navigation. However, these agents typically require elaborate design and expert prompts to solve tasks in specific domains, which limits their adaptability. We introduce AutoManual, a framework enabling LLM agents to autonomously build their understanding through interaction and adapt to new environments. AutoManual categorizes environmental knowledge into diverse rules and optimizes them in an online fashion by two agents: 1) The Planner codes actionable plans based on current rules for interacting with the environment. 2) The Builder updates the rules through a well-structured rule system that facilitates online rule management and essential detail retention. To mitigate hallucinations in managing rules, we introduce case-conditioned prompting strategy for the Builder. Finally, the Formulator agent compiles these rules into a comprehensive manual. The self-generated manual can not only improve the adaptability but also guide the planning of smaller LLMs while being human-readable. Given only one simple demonstration, AutoManual significantly improves task success rates, achieving 97.4% with GPT-4-turbo and 86.2% with GPT-3.5-turbo on ALFWorld benchmark tasks. The source code will be available soon.
@@ 1 Introduction
Several existing methods enable LLM agents to reflect on feedback Reflexion ; AdaPlanner or save successful experiences as skills AdaPlanner ; Voyager ; ExpeL to enhance the performance and reduce the reliance on human-provided examples. However, these reflections and skills have not been well exploited to foster a deeper understanding of the environment. As a result, directly using saved skills as in-context examples can lead to the Path Dependence problem, i.e., the agent blindly replicates the paths of previous successes, failing to appropriately adapt to new scenarios. Such problems are more severe in real-world situations characterized by high variability.
@@
  • We adopt free-form code as the way for the Planner agent to interact with the environment. We introduce a structured rule system that allows the Builder agent to manage multiple types of knowledge from these code-based interactions.
@@ 3.1 AutoManual Overview
Figure 1: AutoMaual Overview. AutoMaual operates in three stages: (1) Building Stage: The Planner agent interacts with the environment by coding actionable plans. After receiving the current trajectory of the Planner, the Builder agent manages rules through the online rule system. (2) Formulating Stage: The Formulator agent formulates the resulting rules into a Markdown manual. (3) Testing Stage: A test-time Planner agent utilizes the manual to complete testing tasks.
@@
Inspired by this online reinforcement learning paradigm, we follow two alternative processes to optimize the rules $\Theta$: 1. The Planner practices the current rules through interaction within an episode. 2. The Builder updates the rules $\Theta$ based on this trajectory. Compared to traditional parameter optimization, sample-inefficient gradient ascent is replaced by text-based rule management. We design a well-structured rule system described in Section 3.3 to ensure the rule updating contributes to rewards. Additionally, to limit the role of human expertise, we only provide a simple example demonstrating the output format to agents. We derive initial rules from this example as the starting point of the optimization.
@@ 3.2 Planner Agent for Interactive Planning
    Free-Form Code: A block of Python code divided into steps. The Planner is encouraged to define helpful functions in the code, which might be reusable in similar scenarios.
@@
Figure 2: The Planner Trajectory. Given the current task and rules, the Planner will interact with the environment through free-form code. Based on the trajectory result, the Planner will generate a corresponding conclusion, which will be saved in the skill or reflection library.
Skill Library and Reflection Library: Apart from rules, we also manage and transmit conclusions from previous episodes, which provide essential details for generating planning code. In the case of Direct or Indirect Success, we save the code block in conclusion as a skill for that task type  [^1] All tasks in ALFWorld are divided into 6 task types, e.g., pick_heat_then_place, look_at_obj_in_light. We only store one skill for each task type, which will be replaced with a new skill. into the skill library AdaPlanner ; Voyager . In the Failure case, we save its conclusion as a reflection for that task type into the reflection library. When a new task comes, the code block of the most similar task is retrieved from the skill library. If there is no existing skill for the new task type, the reflection for that task type will be returned. As mentioned in the Introduction, compared with rules, these skills and reflections contain more programming details but are less generalizable to new scenarios. Considering this Path Dependence problem, we prompt the Planner that the rules should be prioritized.
@@ 3.3 Builder and Consolidator Agents for Rule Management
Figure 3: Case-Conditioned Prompts. Given the current trajectory, the Builder classifies the cause of the major error as "Imperfect Rules" or "Imperfect Agents". Then, the Builder will get the base prompt and corresponding prompt to guide its rule management.
Rule System: We intuitively identify rules as the kinds of knowledge that help task completion, including the analyses of the observed phenomenon $T(o^{\prime}|o,a)$, the mechanism $T(s^{\prime}|s,a)$, and the correlation between the reward $r$ and $\tau_{\rho}$, i.e., the success process or the occurred error. Therefore, unlike ExpeL ExpeL and AutoGuide AutoGuide , which derive general insight from the trajectory, our system categorizes six specific rule types to extract environmental knowledge that targets different aspects of the trajectory. Furthermore, each rule in our system is enhanced with an "Example" attribute to illustrate its application and important details, making it grounded and well-understood. Specifically, each rule in the rule system has these four attributes:
@@
    Rule Type: The type of the rule, options include “Special Phenomenon”, “Special Mechanism”, “Success Process”, “Useful Helper Method”, “Corrected Error” and “Unsolved Error”;
@@ 3.4 Manual Formulation
Once the building stage is complete, we can obtain a set of rules targeted to different situations. Although these rules have been validated through online optimization to ensure their applicability, our next goal is to enhance their readability and global understanding. To achieve this, we introduce the Formulator agent, designed to transform these rules into a user-friendly manual, like a teacher conveys complex subjects through easily digestible lessons. As depicted in Fig 1, the Formulator begins by categorizing all rules based on their target scenarios. This categorization aids in structuring the manual and ensures that related rules are discussed together, which enhances the logical flow and accessibility of the information. For each category, the Formulator drafts an introduction, summarizing the rules it contains and highlighting the key points and overall principles that govern the specific scenarios. Finally, the Formulator compiles the rules and their introductions into a comprehensive manual formatted in Markdown.
@@ 4 Experiments
In line with AdaPlanner AdaPlanner , we conduct the experiments on two interactive environments: (1) ALFWorld ALFWorld is a text-based virtual household environment containing six distinct task types. We run the building stage on 36 tasks (6 tasks for each task type) sampled from the training set of ALFWorld, and each task is run only once. Following previous works Reflexion ; AdaPlanner ; ReAct , we run the testing stage on the validation unseen set containing 134 tasks across these six types. (2) MiniWoB++ MiniWoB is a simulated web environment where agents complete diverse tasks on the Internet by performing keyboard and mouse actions. Prior works RCI ; AdaPlanner selects 9 task types with environmental feedback and 44 task types without feedback from MiniWoB++ tasks. We perform experiments on 9 task types with feedback and all 53 task types. At each stage, we randomly sample 6 tasks for each task type.
During building and formulating stages, we use GPT-4-turbo (gpt-4-1106-preview) as the LLM for all agents. At the testing stage, we equip the Planner agent with GPT-4-turbo or GPT-3.5-turbo (gpt-3.5-turbo-1106), which can evaluate the effect of the generated manual on relatively smaller LLM. More details of the implementation and prompts for AutoManual can be found in the Appendix.
@@
ReAct, Reflexion, and ExpeL provide LLM agents with $12$ human examples ($2$ examples per task type) of ALFWorld. For AdaPlanner, they provide 6 human examples ($1$ example per task type) of ALFWorld as the start of skill discovery. For our methods, agents are provided only one human example of the simplest task (Put) on ALFWorld. On MiniWob++, our agents are provided one human example (search-engine) for tasks with feedback and 4 examples for all tasks. Reflexion agents are allowed to try at most $3$ trials for each task. For AdaPlanner and our methods, we allow the Planner agent to replan at most $3$ times in response to the environmental feedback. To reduce randomness, we performed each experiment three times and reported the average.
@@ 4.1 Main Results
(a) Cross-Task v.s. Single-Task Type.
(b) AutoManual v.s. Skill Library.
@@ Appendix A Limitations
Secondly, the current implementation places all rules directly within the context of LLM, which, while effective in smaller or moderately complex environments, may not scale well to larger, more dynamic settings. For such expansive environments, integrating the rule system with Retrieval-Augmented Generation (RAG) techniques, similar to the approach taken by AutoGuide, could enhance performance by dynamically selecting relevant rules based on the context, thereby managing the cognitive load on the LLM more efficiently.
Lastly, there remains a challenge in ensuring that the Planner agent consistently adheres to the rules outlined in the manual. In practice, the Planner may sometimes ignore these rules and cannot generate the ideal solution code that can be applied to different situations. This indicates a need for additional mechanisms to enforce or verify rule compliance during operations. This issue underscores the potential necessity for developing more sophisticated methods to ensure rule adherence or to introduce more robust validation steps within the planning process.
@@ Appendix C Difference with Prior Methods
2) Online v.s. Offline. ExpeL and AutoGuide extract rules from offline experiences of Reflexion agent. We update rules online by alternating rule practice and rule management. This online rule management can verify the reliability and applicability of the rules in a timely manner, forbidding rules to be armchair general. Additionally, online updated rules can help planners continuously improve their trajectories, making it easier for higher-quality success processes to emerge.
@@ Appendix D Implementation Details
In the building stage of all experiments, the maximum number of rules was set to 12 to balance the context length and rule diversity. We use OpenAI’s Assistant API for all agents to save their history and prevent duplicate inputs. We set all LLMs’ temperatures to 0 and maximum context length to 16000. For AdaPlanner and AutoManual, we allow the Planner agent to replan at most $3$ times on ALFWorld and $6$ times on MiniWob++. In the Building stage on ALFWorld, we use 6 tasks for each task type, a total of 36 tasks, by default. We shuffle all tasks, and when tasks in a task type succeed three times in a row during building, we consider that this task type has been solved and will no longer run it. The API call cost for building and formulating stages is about \$14 in total.

### v4

# AutoManual: Constructing Instruction Manuals by LLM Agents via Interactive Environmental Learning

@@ Abstract
Large Language Models (LLM) based agents have shown promise in autonomously completing tasks across various domains, e.g., robotics, games, and web navigation. However, these agents typically require elaborate design and expert prompts to solve tasks in specific domains, which limits their adaptability. We introduce AutoManual, a framework enabling LLM agents to autonomously build their understanding through interaction and adapt to new environments. AutoManual categorizes environmental knowledge into diverse rules and optimizes them in an online fashion by two agents: 1) The Planner codes actionable plans based on current rules for interacting with the environment. 2) The Builder updates the rules through a well-structured rule system that facilitates online rule management and essential detail retention. To mitigate hallucinations in managing rules, we introduce a case-conditioned prompting strategy for the Builder. Finally, the Formulator agent compiles these rules into a comprehensive manual. The self-generated manual can not only improve the adaptability but also guide the planning of smaller LLMs while being human-readable. Given only one simple demonstration, AutoManual significantly improves task success rates, achieving 97.4% with GPT-4-turbo and 86.2% with GPT-3.5-turbo on ALFWorld benchmark tasks. The code is available at https://github.com/minghchen/automanual.
@@ 1 Introduction
Several existing methods enable LLM agents to reflect on feedback Reflexion ; AdaPlanner or save successful experiences as skills AdaPlanner ; Voyager ; ExpeL to enhance the performance and reduce the reliance on human-provided examples. However, these reflections and skills have not been well exploited to foster a deeper understanding of the environment. As a result, directly using saved skills as in-context examples can lead to the Path Dependence problem, i.e., the agent blindly replicates the paths of previous successes, failing to adapt appropriately to new scenarios. Such problems are more severe in real-world situations characterized by high variability.
@@
  • We adopt actionable code as the way for the Planner agent to interact with the environment. We introduce a structured rule system that allows the Builder agent to manage multiple types of knowledge from these code-based interactions.
@@ 2.2 Self-improvement of LLM Agents
Another series of works PromptAgent ; LLMOptim ; LLMPrompt employs LLM as a prompt optimizer to enhance its own performance. In contrast to our approach, which addresses challenges in unfamiliar environments, these studies focus on enhancing LLM reasoning performance. As a result, their optimized prompts are typically brief and lack environmental knowledge.
@@ 3.1 AutoManual Overview
Figure 1: AutoManual Overview: AutoManual operates in three stages: (1) Building Stage: The Planner agent interacts with the environment by coding actionable plans. After receiving the current trajectory of the Planner, the Builder agent manages rules through the online rule system. (2) Formulating Stage: The Formulator agent formulates the resulting rules into a Markdown manual. (3) Testing Stage: A test-time Planner agent utilizes the manual to complete testing tasks.
@@
Inspired by this online reinforcement learning paradigm, we follow two alternative processes to optimize the rules $\Theta$: 1. The Planner practices the current rules through interaction within an episode. 2. The Builder updates the rules $\Theta$ based on this trajectory. Compared to traditional parameter optimization, sample-inefficient gradient ascent is replaced by text-based rule management. We design a well-structured rule system described in Section 3.3 to ensure the rule updating contributes to rewards. Additionally, to limit the role of human expertise, we only provide a simple example demonstrating the output format to agents. Then, manually derive several initial rules from this example as the starting point of the optimization.
@@ 3.2 Planner Agent for Interactive Planning
    Code: A block of Python code divided into steps. The Planner is encouraged to define helpful functions in the code, which might be reusable in similar scenarios.
@@
Figure 2: The Planner Trajectory: Given the current task and rules, the Planner will interact with the environment through free-form code. Based on the trajectory result, the Planner will generate a corresponding conclusion, which will be saved in the skill or reflection library.
Skill Library and Reflection Library: Apart from rules, we also manage and transmit conclusions from previous episodes, which provide essential details for generating planning code. In the case of Direct or Indirect Success, we save the code block in conclusion as a skill for that task type  [^1] All tasks in ALFWorld are divided into 6 task types, e.g., pick_heat_then_place, look_at_obj_in_light. For each task type, we store only one skill code. into the skill library AdaPlanner ; Voyager . In the Failure case, we save its conclusion as a reflection for that task type into the reflection library. When a new task comes, the code block of the most similar task is retrieved from the skill library. If there is no existing skill for the new task type, the reflection for that task type will be returned. As mentioned in the Introduction, compared with rules, these skills and reflections contain more programming details but are less generalizable to new scenarios, i.e., the Path Dependence problem.
@@ 3.3 Builder and Consolidator Agents for Rule Management
Figure 3: Case-Conditioned Prompts: Given the current trajectory, the Builder classifies the cause of the major error as "Imperfect Rules" or "Imperfect Agents". Then, the Builder will get the base prompt and corresponding prompt to guide its rule management.
Rule System: We intuitively identify rules as the kinds of knowledge that help task completion, including the analyses of the observed phenomenon $T(o^{\prime}|o,a)$, the mechanism $T(s^{\prime}|s,a)$, and the correlation between the reward $r$ and $\tau_{\rho}$, i.e., the success process or the occurred error. Therefore, unlike ExpeL ExpeL and AutoGuide AutoGuide , which derive general insight from the trajectory, our system categorizes six specific rule types to extract environmental knowledge that targets different aspects of the trajectory. Furthermore, each rule in our system is enhanced with Example attribute to illustrate its application and important details, making it grounded and well-understood. Specifically, each rule in the rule system has these four attributes:
@@
    Rule Type: The type of the rule, chosen from [“Special Phenomenon”, “Special Mechanism”, “Useful Helper Method”, “Success Process”, “Corrected Error”, “Unsolved Error”];
@@ 3.4 Manual Formulation
Once the building stage is complete, we can obtain a set of rules targeted to different situations, whose applicability has been validated through online optimization. Our next goal is to enhance their readability and global understanding. To achieve this, we introduce the Formulator agent, designed to transform these rules into a user-friendly manual, analogous to a teacher imparting a wealth of knowledge through easily digestible lessons. As depicted in Fig 1, the Formulator begins by categorizing all rules based on their target scenarios. This categorization aids in structuring the manual and ensures that related rules are discussed together, which enhances the logical flow and accessibility of the information. For each category, the Formulator drafts an introduction, summarizing the rules it contains and highlighting the key points and overall principles that govern the specific scenarios. Finally, the Formulator compiles the rules and their introductions into a comprehensive manual formatted in Markdown.
@@ 4 Experiments
In line with prior works AutoGuide ; AdaPlanner , we conduct the experiments on three interactive environments: (1) ALFWorld ALFWorld is a text-based virtual household environment containing six distinct task types. We run the building stage on 36 tasks (6 tasks for each task type) sampled from the training set of ALFWorld, and each task is run only once. Following previous works Reflexion ; AdaPlanner ; ReAct , we run the testing stage on the validation unseen set containing 134 tasks across these six types. (2) MiniWoB++ MiniWoB is a simulated web environment where agents complete diverse tasks on the Internet by performing keyboard and mouse actions. Prior works RCI ; AdaPlanner selects 9 task types with environmental feedback and 44 task types without feedback from MiniWoB++ tasks. We perform experiments on 9 task types with feedback or on all 53 task types. At each stage, we randomly sample 6 tasks for each task type. (3) WebArena WebArena introduces realistic web environments by emulating the functionality and data of popular websites. This benchmark poses significant challenges for LLM agents due to its large observation and action space, along with tasks that require longer planning horizons. Following AutoGuide AutoGuide , our experiments focus on the Reddit domain within WebArena.
During the building and formulating stages, we use GPT-4-turbo (gpt-4-1106-preview) for all agents. At the testing stage, we equip the Planner agent with GPT-4-turbo or GPT-3.5-turbo (gpt-3.5-turbo-1106), to evaluate the effect of generated manuals on relatively smaller LLM.
@@
ReAct, Reflexion, and ExpeL provide LLM agents with $12$ human examples ($2$ examples per task type) of ALFWorld. For AdaPlanner, they provide 6 human examples ($1$ example per task type) of ALFWorld as the start of skill discovery. For our methods, agents are provided only one human example of the simplest task (Put) on ALFWorld. On MiniWob++, our agents are provided one human example (search-engine) for tasks with feedback and 4 examples for all tasks. On WebArena, our agents are also provided with one human demonstration. To reduce randomness, we performed each experiment three times and reported the average. More details of the implementation and prompts for AutoManual can be found in the Appendix.
@@ 4.1 Main Results
Table 3: Test on WebArena (Reddit).
Methods Examples Suc(%)
ReAct WebArena 2 6.0
AutoGuide AutoGuide 19 43.7
SteP SteP 14 55.0
Planner 1 51.1
AutoManual 1 65.1
(a) Cross-Task v.s. Single-Task Type.
(b) AutoManual v.s. Skill Library.
@@ 5 Conclusion
6 Acknowledgements
Thanks to Dr. Zhe Zeng for her invaluable assistance with the OpenAI API and GPT Plus services. This work was supported in part by The National Nature Science Foundation of China (Grant No: 62273303), in part by Yongjiang Talent Introduction Programme (Grant No: 2022A-240-G).
@@ Appendix A Limitations
Secondly, the current implementation places all rules directly within the context of LLM, which, while effective in smaller or moderately complex environments, may not scale well to larger, more dynamic settings. For such expansive environments, integrating the rule system with Retrieval-Augmented Generation (RAG) techniques, similar to the approach taken by AutoGuide AutoGuide , could enhance performance by dynamically selecting relevant rules based on the context, thereby managing the cognitive load on the LLM more efficiently.
Thirdly, for complex and challenging tasks, the agents of AutoManual are insufficient in exploring the environment, as they only attempt solutions based on current knowledge. To enhance agents’ exploration of unfamiliar environments, it may be necessary to endow the agent with curiosity Voyager or combine it with tree search algorithms ToT .
Lastly, there remains a challenge in ensuring that the Planner agent consistently adheres to the rules outlined in the manual. In practice, the Planner may sometimes ignore these rules (or have hallucinations about observations) and cannot generate the ideal solution code that can be applied to different situations. This indicates a need for additional mechanisms to enforce or verify rule compliance during operations. This issue underscores the potential necessity for developing more sophisticated methods to ensure rule adherence or to introduce more robust validation steps within the planning process.
@@ Appendix C Difference with Prior Methods
2) Online v.s. Offline. ExpeL and AutoGuide extract rules from offline experiences of Reflexion agent. We update rules online by alternating rule practice and rule management. This online rule management can verify the reliability and applicability of the rules in a timely manner, forbidding rules to be armchair general. Additionally, online-update rules can help planners continuously improve their trajectories, making it easier for higher-quality success processes to emerge.
@@ Appendix D Implementation Details
In the building stage of all experiments, the maximum number of rules was set to 12 to balance the context length and rule diversity. We use OpenAI’s Assistant API for all agents to save their history and prevent duplicate inputs. We set all LLMs’ temperatures to 0 and maximum context length to 16000. Reflexion agents are allowed to try at most $3$ trials for each task. For AdaPlanner and AutoManual, we allow the Planner agent to replan at most $3$ times on ALFWorld and $6$ times on MiniWob++ in response to the environmental feedback. In the Building stage on ALFWorld, we use 6 tasks for each task type, a total of 36 tasks, by default. We shuffle all tasks, and when tasks in a task type succeed three times in a row during building, we consider that this task type has been solved and will no longer run it. The API call cost for building and formulating stages is about \$14 in total.